Determine the molarity of 2.50 x 10-6 kg of potassium phosphate in 500.0mL of solution


1
Expert's answer
2021-05-24T04:09:09-0400

n(K3PO4) = m/M = 2.5*10^-3g/212g/mol = 1.18*10^-5 mol

Molarity of K3PO4 = 1.18*10^-5 mol/0.5 L = 2.36*10^-5 M